Operations Report – April 20th

Crew 280 Operations Report 20-04-2023
SOL: 4
Name of person filing report: Cesca Cufí (Crew Engineer) and Neus Sabaté (Mission Specialist)
Non-nominal systems:
Robotic observatory. Rover Spirit.
Notes on non-nominal systems sections:
Replacement parts of Robotic observatory are scheduled to be installed during crew 261 rotation. Corroded battery wire in Rover Spirit. Replacement pending (Sergii). Rover can only be used for short distance EVA.
Spirit rover used: no
Hours: 216.7
Beginning charge: –
Ending charge: –
Currently charging: no
Opportunity rover used: yes
Hours: 121.1
Beginning charge: 100%
Ending charge: 73%
Currently charging: no
Curiosity rover used: no
Hours: 229.8
Beginning charge: –
Ending charge: –
Currently charging: no
Perseverance rover used: yes
Hours: 261.5
Beginning charge: 100%
Ending charge: 56%
Currently charging: no
General notes on rovers:
Perseverance batteries are older and discharge faster. From the station to Cowboy Corner the battery decreased down to 65%.
Summary of Hab operations:
Water use: 23 gal (from 6PM Sol3 to 6PM Sol4). Note: tank refiled yesterday, the current level of water has reached the upper section of the tank, where formula probably gives an overestimation of the water consumption.
Water (static tank): 420 gal (at 6:00PM)
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 18:00 to 21:00 for personal communication and reports.
Summary of suits and radios:
Suits: 3,7,8 and 10.
Radios: EVA radios work nominally. Channel 1 com with the station was lost after passing by Galileo Road, driving North, as expected (from the data provided by Mission support). The EVA-Station com was maintained thanks to the crew’s own radio equipment (Midland HAM radios).
All radio and spacesuit batteries checked before EVA: ok.
Summary of GreenHab operations:
Water use: 12 gal
Heater: off
Supplemental light: off
Harvest: 21 g micro greens, 3 g mint, 116 g red tomatoes, 23 g yellow tomatoes
Summary of Science Dome operations:
Laia Ribas (Green Hab Officer): experiment.
Neus Sabaté (Mission Specialist): experiment.
Dual split: off
Summary of RAM operations: none
Summary of any observatory issues:
Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.
Summary of health and safety issues: none
Summary of other system’s issues:
Hab heater: it turned off about 17PM due to the high winds. Same light code as in Sol3. Crew Engineer troubleshoot the issue with the procedure given by Mission Support the previous sol (Sol3).
Hab outer layer: Hab outer layer is a little loose. When the wind is blowing, the panel is knocking the structure, constantly. The integrity of the structure is not concerned. Replacement of hab outer layer scheduled for October.
Questions, concerns and requests to Mission Support:
Spirit rover: replacement of the corroded wire
Kitchen: knobs replacement
Retrieval of trash in the rear airlock

Copyright © The Mars Society. All rights reserved. | Main Site